A cylindrical canister of external diameter 32 cm and with walls 1 cm thick is filled with water to a height of 50 cm inside canister. The water is then poured into a second cylindrical canister with external diameter 52 cm and with walls 1 cm thick. Assuming that the water does not overfill the canister, to what height does the water fill the second canister?

Answer by Boreal(15235)

The volume of the first cylinder is pi*r^2h, where d=30 cm and r=15 cm, 32-2 for each width.
V=pi*15^2*50=35,342.92 cm^3
The second is 50 cm in diameter or radius of 25 cm.
35342.92=pi*25^2*h
h=18.00 cm